# Via Pietro Giuria 1, I-10125 Turin, Italy

### v1

@@ 2.4 Higher dimensions and non-Abelian theories
The central-plaquette geometry can be generalized to higher dimensions, too. Consider a $s$-simplex with at its boundaries a set of $(s-1)$-simplices, with $s>1$. If the simplex encloses the conical singularity, the size of its boundary is enlarged, since it has to wind around the singularity $n$ times before closing on itself.
@@ 3 Connection with the Hamiltonian formalism
In section 2 we applied the duality transformation to derive the geometry of the replica space for Abelian gauge theories, working in the Lagrangian formalism. Our derivation followed the path depicted in fig. 1, from the lower left corner to the upper right corner, passing through the upper left corner: one starts from a quantum lattice model in the Hamiltonian formalism, where there exists a well-defined notion of Hilbert space, and chooses a factorization $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. For the spin models considered here, such as the Ising model, the factorization can be performed without ambiguities. At this stage one can discretize the path integral that appears in the replica trick, namely
@@
$\sigma^{z}$ and $\sigma^{x}$ are Pauli matrices and, in the boundary term $\sigma_{a}^{z}\sigma_{b}^{z}$, $a$ and $b$ are nearest neighbors. From eq. (37) one can obtain the partition function of a statistical-mechanical system by dividing the (compact) time direction in $N_{\tau}$ intervals of length $a_{\tau}=1/(TN_{\tau})$, Trotterizing the exponential and inserting a resolution of the identity at each discrete time step. If we take, for definiteness, the |$\ket{\sigma^{A}_{r}\sigma^{B}_{r}}$ to be eigenstates of $\sigma^{z}$, the hopping term $\sigma^{z}_{i}\sigma_{j}^{z}$ leads to a nearest-neighbor interaction in the spatial direction, while the $\sigma^{x}$ operators generate interactions in the temporal direction. The replica structure shown in fig. 2 emerges from terms such as
@@
The derivation of the replica space in the presence of superselection sectors discussed in this section is completely general, and can be applied to other spin systems, as well as to higher-dimensional theories. In higher dimensions the explicit derivation of the discretized path integral starting from the Hamiltonian of the theory is less straightforward, and we do not address the calculation here. However, it would be interesting to study how different choices of algebras for gauge theories map to different replica spaces. In particular, the main choices one can make are two: the electric-center algebra and the magnetic center algebra Casini:2013rba ; it seems natural to expect that these two choices correspond to the two different geometries we found in section 2, namely the central-plaquette geometry and the one with the enlarged number of nearest neighbors, while combinations of the two choices of the center might map into combinations of the previous two geometries.
@@ 4 Monte Carlo results for the $\mathbb{Z}_{2}$ gauge theory in three dimensions
where $W_{0\rightarrow 1}$ is the (generalized, dimensionless) work performed to drive a system, initially at equilibrium, from $\lambda=0$ to $\lambda=1$, and the average is performed over an ensemble of different trajectories that the system follows in its out-of-equilibrium evolution. It is important to stress that the equality holds regardless the specific details of how the system is driven off equilibrium: the equality holds irrespectively of the way in which $\lambda$ is varied, as long as this evolution is the same for all trajectories, and that the system is genuinely out of equilibrium for every $\lambda>0$. In recent years, this theorem has been harnessed in high-precision lattice simulations, to study a variety of different quantities Caselle:2016wsw ; Caselle:2018kap ; Francesconi:2020fgi ; Caselle:2022acb ; Caselle:2023uel ; Bonanno:2024udh , including in computations of Rényi entropies through quantum Monte Carlo simulations DEmidio:2019usm ; Zhao:2021njg ; Zhao:2021ghz ; DaLiao:2023pdn ; Pan:2023ysg ; Deng:2023akc ; Song:2023wlg ; DEmidio:2024ggo .
@@
The function fits the data well from large values of $lm_{\mbox{\tiny{g}}}$ down to $lm_{\mbox{\tiny{g}}}\simeq 0.84$. The results of the fit are $A=0.33(3)$, $\alpha=0.360(19)$, with $\chi^{2}_{\tiny\mbox{red}}=0.82$ (where $\chi^{2}_{\tiny\mbox{red}}$ denotes the $\chi^{2}$ divided by the number of degrees of freedom of the fit). Note that the parameter $\alpha$ appearing in eq. (52) controls the exponential decay of the entropic c-function at large distances, and it reduces to $1$ for a free scalar field. For the $\mathbb{Z}_{2}$ gauge theory we obtain a value of $\alpha$ which is smaller than the free-field case, albeit roughly of the same order. A discrepancy with respect to the free case is expected, since the system we are considering is interacting.
Consider now the opposite regime (small $lm_{\mbox{\tiny{g}}}$); the recent ref. Florio:2023mzk reported a study of the two-mode logarithmic negativity in the Schwinger model, which was studied as a measure of entanglement at different length scales. It was found that, at distances shorter than the inverse of the mass of the lightest state in the continuum, the decay of the two-mode negativity is polynomial. Motivated by this observation, we fitted our data for small $lm_{\mbox{\tiny{g}}}$ with a polynomial function
@@
Note that a deviation from the polynomial decay is expected at sufficiently short length scales, since $C_{2}(0)$ is finite. At very short distances, however, the simulations become more expensive in terms of computing resources, since one should get very close to the critical point of the theory and very large lattices are necessary to extrapolate to infinite volume. Despite these difficulties, looking at the data we obtained from finite lattices and for different values of the lattice spacing, as shown in fig. 6, one observes clear convergence towards a finite value for $lm_{\mbox{\tiny{g}}}\rightarrow 0$, even though some data sets appear to tend to a value slightly different from $1$. This is likely a systematic uncertainty due to the finiteness of the lattice spacing and/or of the length of the slab, which become particularly relevant effects in the $lm_{\mbox{\tiny{g}}}\rightarrow 0$ limit.
@@ 5 Conclusions
In order to make contact with Hamiltonian formalism, in section 3 we derived the replica geometry for the $(1+1)$-dimensional Ising model in presence of superselection sectors, finding the same geometry of section 2, with spins on the entangling surface having an enlarged number of nearest neighbors. The derivation is completely general and in principle it might be extended to gauge theories with magnetic- and electric-center algebra.
In the second part of our work, we numerically exploited the duality transformation, simulating the three-dimensional Ising model to compute the entropic c-function associated with the second Rényi entropy of the $\mathbb{Z}_{2}$ gauge theory. Despite its simplicity, this theory displays many phenomena that also characterize $\mathrm{SU}(N)$ gauge theories, such as a mass-gap and confinement. Combining the efficient algorithm Bulgarelli:2023ofi based on Jarzynski’s theorem Jarzynski:1996oqb ; Jarzynski:1997ef and GPU parallelization, we were able to extrapolate our results to the infinite-volume and continuum limits, and to compare our final results with predictions based on the holographic correspondence Klebanov:2007ws .
@@
At short length scales, the entropic c-function scales polynomially with the length of the slab. The exponent of the polynomial decay we obtained fitting our data is compatible with $\frac{1}{2}$ and it might be related to the scaling dimension of some operator in the conformal field theory describing the three-dimensional Ising model at criticality. Note, however that such exponent is likely to be connected in a non-trivial way to the scaling dimension of the twist fields Calabrese:2004eu ; Cardy:2007mb . The transition from polynomial to exponential scaling is observed to occur around a length scale controlled by the mass of the lightest glueball, that is the mass gap of the theory. This behavior resembles the transition that was found in many holographic models dual to a confining background Ryu:2006ef ; Klebanov:2007ws ; Kol:2014nqa and also in Monte Carlo studies of $\mathrm{SU}(N)$ gauge theories Buividovich:2008gq ; Buividovich:2008kq ; Buividovich:2008yv ; Itou:2015cyu ; Rabenstein:2018bri . Moreover, we note that similar results were recently obtained from a study of the two-mode entanglement negativity in the Schwinger model Florio:2023mzk .
In this work we studied the simplest, non-trivial gauge theory which admits a dual spin description. A natural generalization is to study the entropic c-function of other Abelian gauge theories by exploiting the duality transformation. In this respect, a particularly interesting theory is the $\mathrm{U}(1)$ gauge theory in $(2+1)$ dimensions. Moreover, as we previously stated, the same geometries we derived for Abelian theories can in principle be used to study Rényi entropies in non-Abelian gauge theories. We leave these direction of research for future work.
@@ Appendix A Duality transformation and boundary conditions
which reduces to the partition function of the Ising model when $\tau_{i,\mu}=1$ for all $i$ and for all $\mu$. In addition to the global $\mathbb{Z}_{2}$ symmetry, the model defined by eq. (A.1) has a local $\mathbb{Z}_{2}$ symmetry: changing the sign of the gauge fields defined on all links touching an arbitrary site, the partition function of the model is left invariant. For future convenience, note that a generic set of frustrations that is gauge-equivalent to a configuration without any frustrations at all is dual to a set of closed loops in the dual lattice, where the loops are made of links piercing the frustrated links in the direct lattice (see fig. 7).
@@
namely a generic spin-correlator in the $D=2$ Ising model can be expressed in terms of the ratio of the partition function with frustrations piercing paths that connect the various insertions, over the partition function without frustrations.
@@
New solutions to the constraint imposed by the Kronecker delta appearing in eq. (8) can be obtained by choosing any set of closed paths in the lattice dual to the $\mathbb{Z}_{2}$ gauge theory and introducing antiferromagnetic couplings for all plaquettes pierced by those paths.
@@
This is the reason why, in our Monte Carlo study of the entropic c-function of the $\mathbb{Z}_{2}$ gauge theory, we neglected contributions from topologically non-trivial sectors. In addition, this reasoning explains why in section 2 we found that the strong-coupling result expressed by eq. (23) is consistent with the prediction that can be derived in the Hamiltonian formalism, while the weak-coupling one is not. In the weak-coupling regime, the result is spoiled by the effects of non-trivial topology, which are no longer suppressed.
@@ Appendix B Duality in two-dimensional $\mathbb{Z}_{N}$ and $\mathrm{U}(1)$ models
In this section we discuss the duality transformation in systems with $\mathbb{Z}_{N}$ and $\mathrm{U}(1)$ symmetry in a replica space, as a natural generalization of our analysis for the Ising model. A relevant question we want to address here is whether the duality on a replica manifold is affected by the nature of the degrees of freedom of the theory, or just depends on geometric aspects, namely the presence of branch point singularities. Therefore, for these models we will not keep track of the numerical prefactors in the partition function and we will just focus on the geometry of the dual lattice.
@@
where $C_{k}(\beta)$ are the coefficients of the expansion. The partition function then becomes
@@
Consider now a system of $n$ replicas. It is clear that also in this case the duality is only affected by the presence of branch points, therefore we focus our attention on the spins close to one edge of the cut. The geometry is the same as in fig. 3, but with the variables $s$ replaced by $\phi$. We can solve the constraint at site $i$ imposing eq. (B.6),
@@ Appendix D Details of the simulations
In order to convert our results from lattice units to physical units, we set the scale using the study of the Ising model in three dimensions reported in ref. Caselle:1995wn , where the values of $\beta$ at which the deconfinement transition occurs were computed numerically, for a range of temporal extents up to $N_{\tau}=16$. To perform simulations even closer to the critical point and access finer lattices, we extrapolated the results of ref. Caselle:1995wn , using eq. (22) in the same paper, to larger $N_{\tau}$ values. Note that here and in the following we are implicitly trading the $a\to 0$ limit for the equivalent $N_{\tau}\to\infty$ limit.
@@
The fitting function we used for the infinite-volume extrapolation is
@@
The continuum extrapolation was performed in two different ways. The first one consists in approximating the derivative with the backward difference,
@@
which, in principle, has ($\order{a^{2}}$ ) discretization errors. However, with this approach the different points we use for the continuum extrapolations are evaluated at $(l+a/2)m_{\mbox{\tiny{g}}}$, which tends to $lm_{\mbox{\tiny{g}}}$ as $a$ goes to zero but introducing further discretization errors which are linear in $a$. We took the absolute difference between the continuum extrapolations carried out with these two methods as an estimate of (a source of) the systematic error on our results. The final result reported is the value of the mid-point approximation with an uncertainty obtained by summing in quadrature the statistical and the systematic error.

### v2

@@ 2.4 Higher dimensions and non-Abelian theories
The central-plaquette geometry can be generalized to higher dimensions, too. Consider an $s$-simplex with at its boundaries a set of $(s-1)$-simplices, with $s>1$. If the simplex encloses the conical singularity, the size of its boundary is enlarged, since it has to wind around the singularity $n$ times before closing on itself.
@@ 3 Connection with the Hamiltonian formalism
In section 2 we applied the duality transformation to derive the geometry of the replica space for Abelian gauge theories, working in the Lagrangian formalism. Our derivation followed the path depicted in figure 1, from the lower left corner to the upper right corner, passing through the upper left corner: one starts from a quantum lattice model in the Hamiltonian formalism, where there exists a well-defined notion of Hilbert space, and chooses a factorization $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. For the spin models considered here, such as the Ising model, the factorization can be performed without ambiguities. At this stage one can discretize the path integral that appears in the replica trick, namely
@@
$\sigma^{z}$ and $\sigma^{x}$ are Pauli matrices and, in the boundary term $\sigma_{a}^{z}\sigma_{b}^{z}$, $a$ and $b$ are nearest neighbors. From eq. (37) one can obtain the partition function of a statistical-mechanical system by dividing the (compact) time direction in $N_{\tau}$ intervals of length $a_{\tau}=1/(TN_{\tau})$, Trotterizing the exponential and inserting a resolution of the identity at each discrete time step. If we take, for definiteness, the |$\ket{\sigma^{A}_{r}\sigma^{B}_{r}}$ to be eigenstates of $\sigma^{z}$, the hopping term $\sigma^{z}_{i}\sigma_{j}^{z}$ leads to a nearest-neighbor interaction in the spatial direction, while the $\sigma^{x}$ operators generate interactions in the temporal direction. The replica structure shown in figure 2 emerges from terms such as
@@
The derivation of the replica space in the presence of superselection sectors discussed in this section is completely general, and can be applied to other spin systems, as well as to higher-dimensional theories. In higher dimensions the explicit derivation of the discretized path integral starting from the Hamiltonian of the theory is less straightforward, and we do not address the calculation here. However, it would be interesting to study how different choices of algebras for gauge theories map to different replica spaces. In particular, the main choices one can make are two: the electric-center algebra and the magnetic-center algebra Casini:2013rba ; it seems natural to expect that these two choices correspond to the two different geometries we found in section 2, namely the central-plaquette geometry and the one with the enlarged number of nearest neighbors, while combinations of the two choices of the center might map into combinations of the previous two geometries.
@@ 4 Monte Carlo results for the $\mathbb{Z}_{2}$ gauge theory in three dimensions
where $W_{0\rightarrow 1}$ is the (generalized, dimensionless) work performed to drive a system, initially at equilibrium, from $\lambda=0$ to $\lambda=1$, and the average is performed over an ensemble of different trajectories that the system follows in its out-of-equilibrium evolution. It is important to stress that the equality holds regardless of the specific details of how the system is driven off equilibrium: the equality holds irrespectively of the way in which $\lambda$ is varied, as long as this evolution is the same for all trajectories, and that the system is genuinely out of equilibrium for every $\lambda>0$. In recent years, this theorem has been harnessed in high-precision lattice simulations, to study a variety of different quantities Caselle:2016wsw ; Caselle:2018kap ; Francesconi:2020fgi ; Caselle:2022acb ; Caselle:2023uel ; Bonanno:2024udh , including in computations of Rényi entropies through quantum Monte Carlo simulations DEmidio:2019usm ; Zhao:2021njg ; Zhao:2021ghz ; DaLiao:2023pdn ; Pan:2023ysg ; Deng:2023akc ; Song:2023wlg ; DEmidio:2024ggo .
@@
The function fits the data well from large values of $lm_{\mbox{\tiny{g}}}$ down to $lm_{\mbox{\tiny{g}}}\simeq 0.84$. The results of the fit are $A=0.33(3)$, $\alpha=0.360(19)$, with $\chi^{2}_{\tiny\mbox{red}}=0.82$ (where $\chi^{2}_{\tiny\mbox{red}}$ denotes the $\chi^{2}$ divided by the number of degrees of freedom of the fit). Note that the parameter $\alpha$ appearing in eq. (52) controls the exponential decay of the entropic c-function at large distances, and it reduces to $1$ for a free scalar field. For the $\mathbb{Z}_{2}$ gauge theory we obtain a value of $\alpha$ which is smaller than in the free-field case, albeit roughly of the same order. A discrepancy with respect to the free case is expected, since the system we are considering is interacting.
Consider now the opposite regime (small $lm_{\mbox{\tiny{g}}}$); the recent ref. Florio:2023mzk reported a study of the two-mode logarithmic negativity in the Schwinger model, which was studied as a measure of entanglement at different length scales. It was found that, at distances shorter than the inverse of the mass of the lightest state in the continuum, the decay of the two-mode negativity is polynomial. Motivated by this observation, we fitted our data for small $lm_{\mbox{\tiny{g}}}$ with a power-law function
@@
Note that a deviation from the power-law decay is expected at sufficiently short length scales, since $C_{2}(0)$ is finite. At very short distances, however, the simulations become more expensive in terms of computing resources, since one should get very close to the critical point of the theory and very large lattices are necessary to extrapolate to infinite volume. Despite these difficulties, looking at the data we obtained from finite lattices and for different values of the lattice spacing, as shown in figure 6, one observes clear convergence towards a finite value for $lm_{\mbox{\tiny{g}}}\rightarrow 0$, even though some data sets appear to tend to a value slightly different from $1$. This is likely a systematic uncertainty due to the finiteness of the lattice spacing and/or of the length of the slab, which become particularly relevant effects in the $lm_{\mbox{\tiny{g}}}\rightarrow 0$ limit.
@@ 5 Conclusions
In order to make contact with the Hamiltonian formalism, in section 3 we derived the replica geometry for the $(1+1)$-dimensional Ising model in presence of superselection sectors, finding the same geometry of section 2, with spins on the entangling surface having an enlarged number of nearest neighbors. The derivation is completely general and in principle it might be extended to gauge theories with magnetic- and electric-center algebra.
In the second part of our work, we numerically exploited the duality transformation, simulating the three-dimensional Ising model to compute the entropic c-function associated with the second Rényi entropy of the $\mathbb{Z}_{2}$ gauge theory. Despite its simplicity, this theory displays many phenomena that also characterize $\mathrm{SU}(N)$ gauge theories, such as a mass gap and confinement. Combining the efficient algorithm Bulgarelli:2023ofi based on Jarzynski’s theorem Jarzynski:1996oqb ; Jarzynski:1997ef and GPU parallelization, we were able to extrapolate our results to the infinite-volume and continuum limits, and to compare our final results with predictions based on the holographic correspondence Klebanov:2007ws .
@@
At short length scales, the entropic c-function scales as a power of the length of the slab. The exponent of the power-law decay we obtained fitting our data is compatible with $\frac{1}{2}$ and it might be related to the scaling dimension of some operator in the conformal field theory describing the three-dimensional Ising model at criticality. Note, however that such exponent is likely to be connected in a non-trivial way to the scaling dimension of the twist fields Calabrese:2004eu ; Cardy:2007mb . The transition from power-law to exponential scaling is observed to occur around a length scale controlled by the mass of the lightest glueball, that is the mass gap of the theory. This behavior resembles the transition that was found in many holographic models dual to a confining background Ryu:2006ef ; Klebanov:2007ws ; Kol:2014nqa and also in Monte Carlo studies of $\mathrm{SU}(N)$ gauge theories Buividovich:2008gq ; Buividovich:2008kq ; Buividovich:2008yv ; Itou:2015cyu ; Rabenstein:2018bri . Moreover, we note that similar results were recently obtained from a study of the two-mode entanglement negativity in the Schwinger model Florio:2023mzk .
In this work we studied the simplest, non-trivial gauge theory which admits a dual spin description. A natural generalization is to study the entropic c-function of other Abelian gauge theories by exploiting the duality transformation. In this respect, a particularly interesting theory is the $\mathrm{U}(1)$ gauge theory in $2+1$ dimensions. Moreover, as we previously stated, the same geometries we derived for Abelian theories can in principle be used to study Rényi entropies in non-Abelian gauge theories. We leave these directions of research for future work.
@@ Appendix A Duality transformation and boundary conditions
which reduces to the partition function of the Ising model when $\tau_{i,\mu}=1$ for all $i$ and for all $\mu$. In addition to the global $\mathbb{Z}_{2}$ symmetry, the model defined by eq. (A.1) has a local $\mathbb{Z}_{2}$ symmetry: changing the sign of the gauge fields defined on all links touching an arbitrary site, the partition function of the model is left invariant. For future convenience, note that a generic set of frustrations that is gauge-equivalent to a configuration without any frustrations at all is dual to a set of closed loops in the dual lattice, where the loops are made of links piercing the frustrated links in the direct lattice (see figure 7).
@@
namely a generic spin correlator in the $D=2$ Ising model can be expressed in terms of the ratio of the partition function with frustrations piercing paths that connect the various insertions, over the partition function without frustrations.
@@
New solutions to the constraint imposed by the Kronecker delta appearing in eq. (8) can be obtained by choosing any set of closed paths in the lattice dual to the one of the $\mathbb{Z}_{2}$ gauge theory and introducing antiferromagnetic couplings for all plaquettes pierced by those paths.
@@
This is the reason why, in our Monte Carlo study of the entropic c-function of the $\mathbb{Z}_{2}$ gauge theory, we neglected contributions from topologically non-trivial sectors. In addition, the previous arguments explain why in section 2 we found that the strong-coupling result expressed by eq. (23) is consistent with the prediction that can be derived in the Hamiltonian formalism, while the weak-coupling one is not. In the weak-coupling regime, the result is spoiled by the effects of non-trivial topology, which are no longer suppressed.
@@ Appendix B Duality in two-dimensional $\mathbb{Z}_{N}$ and $\mathrm{U}(1)$ models
In this section we discuss the duality transformation in systems with $\mathbb{Z}_{N}$ and $\mathrm{U}(1)$ symmetry in a replica space, as a natural generalization of our analysis for the Ising model. A relevant question we want to address here is whether the duality on a replica manifold is affected by the nature of the degrees of freedom of the theory, or just depends on geometric aspects, namely the presence of branch-point singularities. Therefore, for these models we will not keep track of the numerical prefactors in the partition function and we will just focus on the geometry of the dual lattice.
@@
where the $C_{k}(\beta)$ are the coefficients of the expansion. The partition function then becomes
@@
Consider now a system of $n$ replicas. It is clear that also in this case the duality is only affected by the presence of branch points, therefore we focus our attention on the spins close to one edge of the cut. The geometry is the same as in figure 3, but with the variables $s$ replaced by $\phi$. We can solve the constraint at site $i$ imposing eq. (B.6),
@@ Appendix D Details of the simulations
In order to convert our results from lattice units to physical units, we set the scale using the study of the Ising model in three dimensions reported in ref. Caselle:1995wn , where the values of $\beta$ at which the deconfinement transition occurs were computed numerically, for a range of temporal extents up to $N_{\tau}=16$. To perform simulations even closer to the critical point and access finer lattices, we extrapolated the results of ref. Caselle:1995wn , using eq. (22) in the same paper, to larger $N_{\tau}$ values, as shown in table 1. Note that here and in the following we are implicitly trading the $a\to 0$ limit for the equivalent $N_{\tau}\to\infty$ limit. For additional information on our simulation, see table 2.
@@
The fitting function we used for the infinite-volume extrapolation, shown in figure 8, is
@@
The continuum extrapolation was performed in two different ways, as shown in figure 9. The first one consists in approximating the derivative with the backward difference,
@@
which, in principle, has ($\order{a^{2}}$ ) discretization errors. However, with this approach the different points we use for the continuum extrapolations are evaluated at $(l+a/2)m_{\mbox{\tiny{g}}}$, which tends to $lm_{\mbox{\tiny{g}}}$ as $a$ goes to zero but introducing further discretization errors which are linear in $a$. We took the absolute difference between the continuum extrapolations carried out with these two methods as an estimate of (a source of) the systematic error on our results. The final result reported is the value of the mid-point approximation with an uncertainty obtained by summing in quadrature the statistical and the systematic errors.